Korup National Park (01.01.2016 to 03.31.2016)

You are here

PDF versionPDF version
Year: 
2016
Quarter: 
January - March
Communications activities: 

No communication activity was carried out in our site during this quarter.

Unusual Events at the TEAM Site: 

The effect of el nino is somewhat noticed locally, with higher temperatures. Unfortunately, this is not recorded by the climate station that has been out of order for several months.

New Species: 

None

Protocol Activities: 

Climate station: The climate station is still down (faulty card reader). There is plan for the protocol leader to train a technician locally to download the data directly into the computer.

Trees and Lianas: The data for the fifth recensus of the six 1-ha vegetation plots is completed. Data cleaning is ongoing and uploading will be done as soon as possible.

Terrestrial vertebrates: The pickup of Cameras from Array 1 was done from Jan 12-21, and their deployment to array 2 was done from the February 16-27th. Pickup of the cameras from array 2 was done from March 31st. All photographs of Array 1 were processed and uploaded into deskTEAM. The annotation was not done properly because of some technical issues with the software.

Protocol Problems: 

None

Schedule Problems: 

None

Logistical Problems: 

Two cameras were destroyed by hunters and all the parts including the memry card and batteries were taken away.